OpenStudy (anonymous):

I need to write an equation of the circle with center at (3, 2) and through the point (5, 4). any help would be amazing.

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

where {a,b} is the center and r the radius

OpenStudy (dumbcow):

radius is the distance between (3,2) and (5,4) 

OpenStudy (anonymous):

so this is what I got, am I doing any of this right? (x-3)^2 + (y-2)^2 = 2.8^2

OpenStudy (anonymous):

using distance formula...